Nesmith Corporation's outstanding bonds have a $1,000 par value, a 9% semiannual coupon, 17 years to maturity, and a 12% YTM. What is the bond's price? Round your answer to the nearest cent.
Answer #1

Par Value =1000
Semi annual Coupon =9%*1000/2 =45
Semi annual YTM =12%/2 =6%
Number of Periods =17*2 =34
Price of bond =PV of Coupons+PV of Par Value =45*((1-(1+6%)^-34)/6%)+1000/(1+6%)^34 =784.48